# Information Discovery in e-Commerce

### v1

@@ 4.1.1 Click behavior modeling
where $w$ is the weight vector of the click score. Using stochastic gradient descent [SGD; saad1998online], the authors inferred the likelihood function $p(y{\mid}x,w)$ by applying a sigmoid function over $s(y,x,w)$. Based on these transformed features extracted from gradient boosting decision trees, the authors applied logistic regression as a linear classifier to predict click or no click. Boosted decision trees are able to aggressively reduce the number of active features with only moderate prediction accuracy degradation. The hybrid click model has been widely applied in e-commerce recommendations for candidate ranking (See Section 6.3.2 for more details.).
@@
To explore the feature interactions hidden in data collections, guo2017deepfm proposed a neural network method, i.e., DeepFM, that combines the architectures of factorization machines and deep neural networks. As shown in Fig. 4.3, DeepFM utilizes a wide and deep component to share the same raw input feature vector; this allows the model to learn low- and high-order feature interactions simultaneously. The authors divided DeepFM into the following two components: an FM component and a deep component. These components share the same input features. All parameters are jointly trained for the combined prediction model, as described by Eq. 4.4:
@@
where $\widehat{y}\in(0,1)$ refers to the predicted CTR, $y_{FM}$ is the output of the FM component, and $y_{DNN}$ is the output of the deep component. The authors applied a feed-forward network in the deep component to learn higher-order feature interactions. More DeepFM model-based deep learning methods have been proposed to address the CTR prediction problem, including deep convolutional neural networks (CNNs) [chan2018convolutional] and deep interest neural networks [zhou2018deep, feng2019deep, li2019multi, zhou2019deep, chen2021deepctr, zhu2022tkdectr, guo2022icdectr, cheng2022dynamic]. All of the above-mentioned deep neural networks have significantly contributed to the optimization of item ranking in e-commerce searches and recommendations; this will be further discussed in Sections 5.4 and  6.3.3, respectively.
@@
Additionally, to ensure consistent evaluation and comparison of CTR prediction models, benchmark frameworks such as the open benchmarking for CTR [zhu2021open] and BARS-CTR [zhu2022bars] have been introduced. These frameworks provide a standardized way to evaluate model performance across different datasets, improving reproducibility and promoting further advancements in CTR research.
@@ 4.1.2 Post-click behavior tracking
As we have discussed in Section 3.2.2, post-click behavior plays an important role in modeling for e-commerce users in search (Section 5) and recommendation scenarios (Section 6). In the literature, multiple studies have focused on applying various types of interaction signals to model post-click behaviors in search and recommendation scenarios. sculley2009predicting measured users’ post-click experience by evaluating the corresponding bounce rate. The model proposed by zhong2010incorporating leverages both user clicks on the search page and post-clicks beyond the search page to provide an unbiased estimation of document relevance. lalmas2015promoting investigated how viewport time can be used to measure user attention level as an engagement metric. o2016leveraging leveraged user interactions as signals within the clicked items to enhance the search results. wan2018item determined the monotonic dependency between explicit user signals and more implicit signals to improve recommender systems. lu2018between proposed a preference prediction model to predict user actual preferences for the clicked items by taking into account multiple post-click interactions.
@@
Dwell time. As we have discussed in Section 3.2.2, dwell time is the most common evaluation metric for the analysis of post-click user behavior [yin2013silence]. Accordingly, yin2013silence built a graphical model that focuses on utilizing explicit user feedback and dwell time to predict user preferences in e-commerce recommendations. yi2014beyond showed that integrating dwell time into the learning objective or learning weight results in better recommendation performance than pure predictions of the CTR. rosales2012post, chapelle2015simple utilized dwell time as a proxy of post-click experience in online advertising to improve the ranking performance. bogina2017incorporating explored the value of incorporating dwell time for session-based recommendations by boosting items above the preassigned dwell time threshold. Modeling user behavior by taking into account dwell time has been demonstrated to facilitate e-commerce recommendation performance. In Section 6.2.3, we will discuss more studies that focused on modeling sequential user dynamics by using dwell time.
@@
User return modeling. There is a limited amount of work on modeling user returns in e-commerce, especially when user returns depend heavily on the quality of the provided service [lo2016understanding, zhouwsdm2018]. The model developed by zhouwsdm2018 provides rich user interfaces after a user clicks an item. For instance, it encourages users to visit different item modules or sub-pages, i.e., to read comments or click on pictures embedded within the item page; this generates a large amount of heterogeneous post-click behavior. Three problems pose a challenge for attempts to model micro-behavior on e-commerce platforms: (i) Sparseness and high dimensionality of the user representation; (ii) Sequential information of micro-behavior; and (iii) Diverse effects of micro-behavior. To address these three challenges, zhouwsdm2018 proposed the framework shown in Fig. 4.4, which consists of five layers: an input layer, an embedding layer to solve the problems of sparseness and high dimensionality, an RNN layer to model sequential information, an attention layer to capture the diverse effects of micro-behavior, and an output layer.
@@
where $p_{v}\in\mathbb{R}^{V}$ is a one-hot indicator vector where $p_{v}(i)=1$ if $x_{i}$ is about the $i$-th product and other entities are zero. Similarly, $a_{m}\in\mathbb{R}^{M}$ and $d_{k}\in\mathbb{R}^{K}$ are indicator vectors for activities and dwell time, respectively. Each indicates a unique element in the product set $P$, activity set $A$, and dwell time set $D$, respectively. Here, the vocabulary sizes of $P$, $A$, and $D$ are $V$, $M$, and $K$, respectively, and there are $V\times M\times K$ tuples in total. To address sparseness and high-dimensionality problems, the authors designed an embedding layer to transform the input $x_{t}$ into a low-dimensional dense vector $e_{t}$, which is formally defined as
@@
where $W_{P}\in R^{d_{P}\times V}$, $W_{A}\in R^{d_{A}\times M}$, and $W_{D}\in R^{d_{D}\times K}$, where $d_{P}\ll N$, $d_{A}\ll M$, and $d_{D}\ll K$ are the number of latent dimensions for products, activities, and dwell time, respectively. The initial weights of $W_{P}$, $W_{A}$, and $W_{D}$ are trained by applying word2vec [mikolov2013efficient]. Additionally, the final embedding of $x_{t}$ is the concatenation of three embeddings. To capture the sequential information of micro-behavior, the authors built an RNN layer. The output of the embedding layer $e_{t}$ is the input of the RNN layer. The $t$-th hidden state unit output is calculated as
@@
where $\sigma(\cdot)$ is a non-linear activation function, e.g., ReLU, sigmoid, or tanh; $W_{eh}\in R^{d_{h}\times d_{e}}$, $W_{hh}\in R^{d_{h}\times d_{h}}$, and $b_{i}\in R^{d_{h}}$. To capture the effects of micro-behavior, the authors decided to introduce an attention layer [attention1] that assigns proper weights to each hidden unit; this helps to obtain a more balanced output. The attention weight is mapped from the hidden layer vector to a real valued score by the function $\sigma(\cdot)$. To achieve sufficient expressive ability, the function $\sigma(\cdot)$ is typically implemented by a neural network layer. Then, the final output is an attention weighted pooling of the RNN layer. To exploit the different transition patterns between items and operations in micro-behavior modeling, meng2020sigirclick incorporated the item knowledge into a joint user modeling framework including a recurrent neural network and a graph neural network. To incorporate the micro-behavior information in the iterative process of user behavior modeling, jianhao2022session modeled a user session as a fine-grained sequence of micro-behaviors and proposed a self-attention mechanism to encode the dyadic relations of micro-behaviors.
Experimental results have confirmed that post-click user modeling can provide deeper insight into user behavior, which is utilized to advance e-commerce search and recommender systems by successfully modeling the sequential dynamics in the candidate retrieval stage [zhouwsdm2018]. However, post-click behaviors are often sparse in real-world scenarios, making it challenging to supplement large-scale implicit feedback. To address this, recent studies have integrated post-clicks with other user behaviors. wen2019leveraging proposed a generic probabilistic framework to fuse click and post-click feedback in recommender systems. wang2021clicks revealed the importance of mitigating the clickbait issue from click behaviors, and applied causal inference to establish a causal graph to reformulate the process. In Section 6.2.3, we discuss further studies that have focused on integrating post-click tracking into e-commerce recommendation.
@@ 4.1.3 Purchase-intent modeling
As it is widely applied in search and recommendation approaches, purchase-intent prediction is another important task in e-commerce modeling [qiu2015predicting, kooti2016portrait, lo2016understanding, wan2017modeling]. According to bellman1999predictors, the volume of the online activity of a customer proves useful when predicting the occurrence of a future purchase. Statistical models of customer purchase behavior have been studied for decades. Early research on purchase behavior modeling was based on statistical approaches, e.g., negative binomial distribution models [bearden1999handbook, kooti2016portrait]. Features related to information gathering and the purchase potential (e.g., monetary resources and product values) also help to predict the purchase intention [bearden1999handbook, hansen2004predicting, pavlou2006understanding].
@@
Feature-based methods. User-aware features have been successfully applied to predict purchase intent in e-commerce scenarios. qiu2015predicting proposed a pipeline-based purchase-prediction approach that includes three main components. First, the authors utilized associations between products to predict the needs of customers; then, they combined collaborative filtering and a hierarchical Bayesian discrete choice model enable customer preference learning; lastly, they constructed a support vector regression-based model to calculate the popularity of products. After analyzing user behavior on Pinterest, lo2016understanding proposed a predictor to detect a user’s purchase intent. The authors applied five kinds of features: demographics, activity, action-type, content, and temporal features. wan2017modeling also proposed a three-stage model to predict the purchase behavior on a real-world e-commerce portal. To identify who can be converted to regular loyal buyers and then targeted to reduce promotion cost, liu2016repeat proposed a solution for repeat buyer prediction; they collected a large number of features to capture the preferences and behavior of users, characteristics of merchants, brands, categories, and items, and the interactions among them. hendriksen-2020-analyzing analyzed the potential of long-term historical records (from logged-in users) to more accurately and reliably predict purchase intent. Additionally, ariannezhad-2021-understanding showed that data that provides information on customer behavior in one channel (e.g., online) can help to predict purchase intent in other channels (e.g., offline). Social media has become another important source of information to help explore consumer purchase intentions [mishne-2006-deriving, zhang2013predicting, ding2015mining]. zhang2013predicting explored whether users’ social media information is correlated with their e-commerce profiling categories. Accordingly, the authors leveraged correlations to build machine learning algorithms to predict user purchase behavior.
@@
CVR prediction methods. Conversion rate (CVR) prediction is another way to predict the user purchase intention on e-commerce platforms. CVR calculates the proportion of users who will eventually convert after clicking [lee2012estimating, yang2016large, lu2017practical, wen2019multi, su2020attention, wen2020entire, yasui2020feedback, yang2021capturing, hou2021conversion, li2021conversionpred]. Because conversions are extremely rare, CVR modeling is very challenging. CVR can be split into the following two categories: post-view conversion and post-click conversion, i.e., conversion after viewing an item without having clicked it, and conversion after having clicked the item, respectively. Most approaches focus on the task of post-click conversion. wen2019multi proposed a decision tree ensemble model, i.e., ldcTree, that exploits deep cascade structures and applies cross-entropy based feature representations. Nonetheless, there are still $3$ challenges in CVR estimation: data sparsity, sample selection bias, and delayed feedback. The data sparsity problem reflects the insufficiency of click samples in training data. Sample selection bias refers to the systematic difference in the data distribution between the training space and inference space [wen2020entire]. Fig. 4.5 illustrates the sample selection bias problem related to the development of an efficient industrial-level recommender system.
@@
To address the data sparsity problem, ma2018entire proposed the entire space multi-task model, which aims to apply multi-task learning to accomplish two subtasks of predicting the post-view click-through rate and post-view conversion rate. wen2020entire observed that users always engaged in abundant purchase-related actions after clicking. Thus, they proposed a deep neural network method within a multi-task learning framework to decompose post-click behavior to predict CVR. The authors were able to distinguish between purchase-related actions and other actions, which, taken together, can be used to form a probabilistic sequential user behavior graph.
All the above multi-task strategies are also helpful in alleviating the selection bias problem. Furthermore, zhang2020large proposed a doubly robust estimation method to debias CVR prediction. yasui2020feedback proposed a dual learning method to simultaneously address the delayed feedback problem and the selection bias problem. To reduce the variance of doubly robust loss to enhance model robustness, guo2021enhanced enhanced a more robust doubly robust approach for debiasing post-click conversion rate estimation. But the authors didn’t directly control the bias and the variance in an effective way. To address this problem, dai2022kddcvr proposed a generalized framework of doubly robust learning, which unifies the existing doubly robust methods. Based on this framework, two new doubly robust methods were proposed to control the bias and mean squared error.
Starting to address the delayed feedback problem, chapelle2014modeling proposed the delayed feedback model to optimize CVR as a joint probability over the predicted CVR and the delayed time distribution. yoshikawa2018nonparametric extended the delayed feedback model to a non-parametric model. su2020attention focused on post-click calibration in CVR modeling. The authors extracted pre-trained embeddings from impressions/clicks to enhance the conversion models; they proposed an inner/self-attention mechanism to capture the fine-grained personalized product purchase interests. To estimate unbiased CVR in the online settings, yang2021capturing proposed the elapsed-time sampling delayed feedback model to track relations between the observed conversion distribution and the true conversion distribution. li2021conversionpred utilized an idealized dataset for training a prophet model that can use the data properly, and then learned the actual model by imitating the prophet. chen2022kddasymptotically confirmed the importance of dividing observed samples in a more granular manner, and hence proposed an unbiased importance sampling method with two-step optimization to address the delayed feedback issue.

### v3

@@ 4.1.1 Click behavior modeling
where $w$ is the weight vector of the click score. Using stochastic gradient descent [SGD; saad1998online], the authors inferred the likelihood function $p(y{\mid}x,w)$ by applying a sigmoid function over $s(y,x,w)$. Based on these transformed features, the authors applied logistic regression to predict a click or no click. Boosted decision trees are able to aggressively reduce the number of active features with only moderate prediction accuracy degradation. The hybrid click model is widely applied in e-commerce recommendations for candidate ranking (see Section 6.3.2 for more details.).
@@
To explore the feature interactions hidden in data collections, guo2017deepfm propose a neural network method, i.e., DeepFM, that combines the architectures of factorization machines and deep neural networks. As shown in Fig. 4.3, DeepFM uses a wide and deep component to share the same raw input feature vector; this allows the model to learn low- and high-order feature interactions simultaneously. All parameters are jointly trained for the combined prediction model, as described by Eq. 4.4:
@@
where $\widehat{y}\in(0,1)$ refers to the predicted CTR, $y_{FM}$ is the output of the FM component, and $y_{DNN}$ is the output of the deep component. The authors apply a feed-forward network in the deep component to learn higher-order feature interactions.
More DeepFM model-based deep learning methods have been proposed to address the CTR prediction problem, including deep convolutional neural networks (CNNs) [chan2018convolutional] and deep interest neural networks [zhou2018deep, feng2019deep, li2019multi, zhou2019deep, chen2021deepctr, zhu2022tkdectr, guo2022icdectr, cheng2022dynamic]. All of the above-mentioned deep neural networks have significantly contributed to the optimization of item ranking in e-commerce search and recommendation; this will be further discussed in Sections 5.4 and  6.3.3, respectively.
@@
To ensure consistent evaluation and comparison of CTR prediction models, benchmark frameworks such as the open benchmarking for CTR [zhu2021open] and BARS-CTR [zhu2022bars] have been introduced. These frameworks provide a standardized way to evaluate model performance across different datasets, improving reproducibility and promoting further advancements in CTR prediction research.
@@ 4.1.2 Post-click behavior tracking
As we have discussed in Section 3.2.2, post-click behavior plays an important role in modeling for e-commerce users in search (Section 5) and recommendation scenarios (Section 6). Multiple studies have focused on applying various types of interaction signals to model post-click behaviors in search and recommendation scenarios. sculley2009predicting measure users’ post-click experience by evaluating the corresponding bounce rate. The model proposed by zhong2010incorporating uses both user clicks on the search page and post-clicks beyond the search page to provide an unbiased estimation of document relevance. lalmas2015promoting investigate how viewport time can be used to measure user attention level as an engagement metric. o2016leveraging use user interactions as signals within the clicked items to enhance the search results. wan2018item determine the monotonic dependency between explicit user signals and more implicit signals to improve recommender systems. lu2018between propose a preference prediction model to predict user actual preferences for the clicked items by taking into account multiple post-click interactions.
@@
Dwell time. As we have discussed in Section 3.2.2, dwell time is the most common evaluation metric for the analysis of post-click user behavior [yin2013silence]. Accordingly, yin2013silence built a graphical model that focuses on using explicit user feedback and dwell time to predict user preferences in e-commerce recommendations. yi2014beyond show that integrating dwell time into the learning objective or learning weight results in better recommendation performance than pure predictions of the CTR. rosales2012post, chapelle2015simple use dwell time as a proxy of post-click experience in online advertising to improve the ranking performance. bogina2017incorporating explore the value of incorporating dwell time for session-based recommendations by boosting items above the preassigned dwell time threshold. Modeling user behavior by taking into account dwell time has been shown to facilitate e-commerce recommendation performance. In Section 6.2.3, we will discuss more studies that focused on modeling sequential user dynamics by using dwell time.
@@
User return modeling. There is a limited amount of work on modeling user returns in e-commerce, especially when user returns depend heavily on the quality of the provided service [lo2016understanding, zhouwsdm2018]. The model developed by zhouwsdm2018 provides rich user interfaces after a user clicks an item. For instance, it encourages users to visit different item modules or sub-pages, i.e., to read comments or click on pictures embedded within the item page; this generates a large amount of heterogeneous post-click behavior. Three problems pose a challenge for attempts to model micro-behavior on e-commerce platforms: (i) Sparseness and high dimensionality of the user representation; (ii) Sequential information of micro-behavior; and (iii) Diverse effects of micro-behavior. To address these three challenges, zhouwsdm2018 propose the framework shown in Fig. 4.4, which consists of five layers: an input layer, an embedding layer to solve the problems of sparseness and high dimensionality, an RNN layer to model sequential information, an attention layer to capture the diverse effects of micro-behavior, and an output layer.
@@
where $p_{v}\in\mathbb{R}^{V}$ is a one-hot indicator vector where $p_{v}(i)=1$ if $x_{i}$ is about the $i$-th product and other entities are zero. Similarly, $a_{m}\in\mathbb{R}^{M}$ and $d_{k}\in\mathbb{R}^{K}$ are indicator vectors for activities and dwell time, respectively. Each indicates a unique element in the product set $P$, activity set $A$, and dwell time set $D$, respectively. Here, the vocabulary sizes of $P$, $A$, and $D$ are $V$, $M$, and $K$, respectively, and there are $V\times M\times K$ tuples in total. To address sparseness and high-dimensionality problems, the authors design an embedding layer to transform the input $x_{t}$ into a low-dimensional dense vector $e_{t}$:
@@
where $W_{P}\in R^{d_{P}\times V}$, $W_{A}\in R^{d_{A}\times M}$, and $W_{D}\in R^{d_{D}\times K}$, where $d_{P}\ll N$, $d_{A}\ll M$, and $d_{D}\ll K$ are the number of latent dimensions for products, activities, and dwell time, respectively. The initial weights of $W_{P}$, $W_{A}$, and $W_{D}$ are trained by applying word2vec [mikolov2013efficient]. Additionally, the final embedding of $x_{t}$ is the concatenation of three embeddings. To capture the sequential information of micro-behavior, the authors build an RNN layer. The output of the embedding layer $e_{t}$ is the input of the RNN layer. The $t$-th hidden state unit output is calculated as
@@
where $\sigma(\cdot)$ is a non-linear activation function, e.g., ReLU, sigmoid, or tanh; $W_{eh}\in R^{d_{h}\times d_{e}}$, $W_{hh}\in R^{d_{h}\times d_{h}}$, and $b_{i}\in R^{d_{h}}$. To capture the effects of micro-behavior, the authors introduce an attention layer [attention1] that assigns proper weights to each hidden unit; this helps to obtain a more balanced output. The attention weight is mapped from the hidden layer vector to a real valued score by the function $\sigma(\cdot)$. To achieve sufficient expressive ability, the function $\sigma(\cdot)$ is typically implemented by a neural network layer. Then, the final output is an attention weighted pooling of the RNN layer. To exploit the different transition patterns between items and operations in micro-behavior modeling, meng2020sigirclick incorporate item knowledge into a joint user modeling framework including a recurrent neural network and a graph neural network. To incorporate the micro-behavior information in the iterative process of user behavior modeling, jianhao2022session model a user session as a fine-grained sequence of micro-behaviors and proposed a self-attention mechanism to encode the dyadic relations of micro-behaviors.
Experimental results have confirmed that post-click user modeling can provide deeper insights into user behavior, which is used to advance e-commerce search and recommender systems by successfully modeling the sequential dynamics in the candidate retrieval stage [zhouwsdm2018]. However, post-click behaviors are often sparse in real-world scenarios, making it challenging to supplement large-scale implicit feedback. To address this, recent studies have integrated post-clicks with other user behaviors. wen2019leveraging describe a generic probabilistic framework to fuse click and post-click feedback in recommender systems. wang2021clicks reveal the importance of mitigating the clickbait issue from click behaviors, and apply causal inference to establish a causal graph to reformulate the process. In Section 6.2.3, we discuss further studies that have focused on integrating post-click tracking into e-commerce recommendation.
@@ 4.1.3 Purchase-intent modeling
Purchase-intent prediction is another important task in e-commerce modeling [qiu2015predicting, kooti2016portrait, lo2016understanding, wan2017modeling]. According to bellman1999predictors, the volume of online activities of a customer proves useful when predicting the occurrence of a future purchase. Statistical models of customer purchase behavior have been studied for decades. Early research on purchase behavior modeling was based on statistical approaches, e.g., negative binomial distribution models [bearden1999handbook, kooti2016portrait]. Features related to information gathering and the purchase potential (e.g., monetary resources and product values) also help to predict the purchase intention [bearden1999handbook, hansen2004predicting, pavlou2006understanding].
@@
Feature-based methods. User-aware features have been successfully applied to predict purchase intent in e-commerce scenarios. qiu2015predicting put forward a pipeline-based purchase-prediction approach that includes three main components. First, the authors use associations between products to predict the needs of customers; then, they combine collaborative filtering and a hierarchical Bayesian discrete choice model enable customer preference learning; lastly, they construct a support vector regression-based model to calculate the popularity of products. After analyzing user behavior on Pinterest, lo2016understanding propose a predictor to detect a user’s purchase intent. The authors apply five kinds of feature: demographics, activity, action-type, content, and temporal features. wan2017modeling also introduce a three-stage model to predict the purchase behavior on a real-world e-commerce portal. To identify who can be converted to regular loyal buyers and then targeted to reduce promotion cost, liu2016repeat describe a solution for repeat buyer prediction; they collect a large number of features to capture the preferences and behavior of users, characteristics of merchants, brands, categories, and items, and the interactions among them. hendriksen-2020-analyzing analyze the potential of long-term historical records (from logged-in users) to more accurately and reliably predict purchase intent. Additionally, ariannezhad-2021-understanding show that data that provides information on customer behavior in one channel (e.g., online) can help to predict purchase intent in other channels (e.g., offline). Social media has become another important source of information to help explore consumer purchase intentions [mishne-2006-deriving, zhang2013predicting, ding2015mining]. zhang2013predicting explore whether users’ social media information is correlated with their e-commerce profiling categories. Accordingly, the authors use correlations to build machine learning algorithms to predict user purchase behavior.
@@
CVR prediction methods. Conversion rate (CVR) prediction is another way to predict the user purchase intention on e-commerce platforms. CVR calculates the proportion of users who will eventually convert after clicking [lee2012estimating, yang2016large, lu2017practical, wen2019multi, su2020attention, wen2020entire, yasui2020feedback, yang2021capturing, hou2021conversion, li2021conversionpred]. Because conversions are extremely rare, CVR modeling is very challenging. CVR can be split into the following two categories: post-view conversion and post-click conversion, i.e., conversion after viewing an item without having clicked it, and conversion after having clicked the item, respectively. Most approaches focus on the task of post-click conversion. wen2019multi propose a decision tree ensemble model, i.e., ldcTree, that exploits deep cascade structures and applies cross-entropy based feature representations. Nonetheless, there are still three challenges in CVR estimation: data sparsity, sample selection bias, and delayed feedback. The data sparsity problem reflects the insufficiency of click samples in training data. Sample selection bias refers to the systematic difference in the data distribution between the training space and inference space [wen2020entire]. Fig. 4.5 illustrates the sample selection bias problem related to the development of an efficient industrial-level recommender system.
@@
To address the data sparsity problem, ma2018entire consider the entire space multi-task model, which aims to apply multi-task learning to accomplish two subtasks of predicting the post-view click-through rate and post-view conversion rate. wen2020entire observe that users always engaged in abundant purchase-related actions after clicking. Thus, they propose a deep neural network method within a multi-task learning framework to decompose post-click behavior to predict CVR. The authors distinguish between purchase-related actions and other actions, which, taken together, can be used to form a probabilistic sequential user behavior graph.
The above multi-task strategies are also helpful in alleviating the selection bias problem. Furthermore, zhang2020large detail a doubly robust estimation method to debias CVR prediction. yasui2020feedback provide a dual learning method to simultaneously address the delayed feedback problem and the selection bias problem. To reduce the variance of doubly robust loss to enhance model robustness, guo2021enhanced enhance a more robust doubly robust approach for debiasing post-click conversion rate estimation. But the authors do not directly control the bias and the variance in an effective way. To address this problem, dai2022kddcvr propose a generalized framework of doubly robust learning, which unifies the existing doubly robust methods. Based on this framework, two new doubly robust methods were proposed to control the bias and mean squared error.
chapelle2014modeling describe a delayed feedback model to optimize CVR as a joint probability over the predicted CVR and the delayed time distribution. yoshikawa2018nonparametric extend this delayed feedback model to a non-parametric model. su2020attention focus on post-click calibration in CVR modeling. The authors extract pre-trained embeddings from impressions/clicks to enhance the conversion models; they propose an inner/self-attention mechanism to capture the fine-grained personalized product purchase interests. To estimate unbiased CVR in the online settings, yang2021capturing introduce a elapsed-time sampling delayed feedback model to track relations between the observed conversion distribution and the true conversion distribution. li2021conversionpred use an idealized dataset for training a prophet model that can use the data properly, and then learn the actual model by imitating the prophet. chen2022kddasymptotically confirm the importance of dividing observed samples in a more granular manner, and hence propose an unbiased importance sampling method with two-step optimization to address the delayed feedback issue.